For a SKF Single-Row 02-80 mm Deep-Groove ball bearing, an axial load Fa of 7650 N and a radial load Fr of 15 kN are applied with the

inner ring stationary. Estimate the bearing life at a speed of 740

rev/min with a reliability of 99.99%. SKF rates its bearings for 1 million

revolutions. The Weibull parameters are, x0 = 0.02, θ = 4.459, and b =

1.483


1
Expert's answer
2021-07-12T03:36:56-0400

given Data:


Rotational speed "n = 740 \frac{rev}{min}"


Dynamic equivalent radial load


"P_r = XF_r + YF_a =0.02\times 15000 +1.483\times 7650 =11644.95 N"


"C_r = 7650 N"

Bearing service life (L10h) is calculated

"L_{10h}=\frac{10^6}{n}(\frac{C_r}{P_r})^{3}"


"L_{10h}=\frac{10^6}{740}(\frac{11644.95}{7650})^{3} = 4766.47" hr